OpenStudy (dumptybaby):

If I solve a system and get infinite many solutions, will it be linearly independent?

OpenStudy (irishboy123):

other way round:(

OpenStudy (dumptybaby):

So for it to be linearly independent it has to only have trivial solutions (nothing else) right?

OpenStudy (irishboy123):

I wouldn't normally comment on stuff like this because I have a working rather than theoretical knowledge of it,......., but you do seem to be in bit of a pickle :( Summary: For nxn matrix A, and nx1 vectors \(\mathbf x\) and \(\mathbf b\) in system \(A \mathbf x = \mathbf b\), then if A has a determinant \(A^{-1}\), it follows that there is a non trivial solution: \(A^{-1} A \mathbf x = A^{-1} \mathbf b , \implies \mathbf x = A^{-1} \mathbf b \) A hole bunch of other stuff follows too, especially the fact that the column vectors of A will be **linearly independent** and the row vectors of A will be **linearly independent**. And that might be what you are interested in, maybe. If A is non-invertible, then yes there is something wrong with your system. When the determinant is zero you can an infinite number of, or no solutions. Here is a simple and pretty crap 2x2 example. In the first, you are solving \(y = 2x + 1, y = 2x + 1\), infinite solutions: \(\left( \begin{matrix} 2 & -1 \\ 2 & -1 \end{matrix} \right) \left( \begin{matrix} x \\ y \end{matrix} \right) = \left( \begin{matrix} -1 \\ -1 \end{matrix} \right)\) In the second, you are solving \(y = 2x + 1, y = 2x + 2\), no solutions: \(\left( \begin{matrix} 2 & -1 \\ 2 & -1 \end{matrix} \right) \left( \begin{matrix} x \\ y \end{matrix} \right) = \left( \begin{matrix} -1 \\ -2 \end{matrix} \right)\) The determinant tells you something is wrong. And look at the row and column vectors. All that said, I am more familiar with **homogeneous** systems, \(A \mathbf x = \mathbf 0 \), where you may want there to be no inverse as in the case of systems of linear differential equations. Otherwise you could say the same thing, namely that \( \mathbf x = A^{-1} \mathbf 0 \implies \mathbf x= \mathbf O\) and you're stuffed with a trivial solution. This is where eigenvalues kick in. They avoid trivial solutions but denying the matrix an inverse.

OpenStudy (irishboy123):

PS This is the back page of Gilbert Strang's book. He presents this as Lin Alg for nxn matrices "In a Nutshell". And if he thinks that, it is worth taking seriously :)  His online MIT course is legendary but I doubt you have the time to watch much of it as he goes through in splendid detail. and the URL is case that doesn't show up too well https://i.gyazo.com/a436f7c20d713c1a61b2f10a11ff4c36.png Good luck